OpenStudy (anonymous):

I need to do a proof on If n is even, then n^2 is even.

OpenStudy (psymon):

Well, any number that is even can be written in the form of 2k. Because 2 times any number will be even. Therefore if n = 2k and 2k is even, then n^2 = (2k)^2. (2k)^2 = 4k^2, which is the same as 2(2k^2). So because we can write n as 2k, an even number, and n^2 can be written as 2(2k^2), then we can say that n^2 msut also be even.

OpenStudy (anonymous):

that doesnt actually prove n^2 is even. to prove this you need to show n^2 is divisible by 2 or n^2=2k. then show n=something that will equal 2k

OpenStudy (anonymous):

when squared

OpenStudy (psymon):

If you know how and know im wrong then sounds like you wouldnt need our help.

OpenStudy (psymon):

And saying n = 2k, i would think, would be more valid because integers arent closed under division.